OpenStudy (anonymous):

Factor x^2 – 5x + 24. A. (x – 12)(x – 2) B. (x – 3)(x – 8) C. (x – 4)(x – 6) D. prime

OpenStudy (paki):

x2-8x+3x+24... solve now....

OpenStudy (shiraz14):

It should be x^2 - 11x + 24

OpenStudy (anonymous):

Thats how they gave me the question..

OpenStudy (shiraz14):

Then the answer is D.

OpenStudy (anonymous):

Thanks

OpenStudy (shiraz14):

You're welcome

OpenStudy (anonymous):

Do you know ac method to find factor s ?

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

Okay first compare this equation with standard quardratic equation ax^2 + bx+c =0 so coan you tell me a = ----? , b = -----? , c = -------?

OpenStudy (anonymous):

C is 24?

OpenStudy (anonymous):

B is 5?

OpenStudy (anonymous):

I dont know what A is..

OpenStudy (paki):

A is 1

OpenStudy (anonymous):

Oh ok, I can see that.

OpenStudy (paki):

i have solved this question by factors... but the answer is not matching with your one by one (negative) sign....

OpenStudy (anonymous):

By ac method a*c = 24 so we find such two numbers whose product is 24 & sum is -5x so we will get x^2 - 8x + 3x + 24 = 0 now factor out this x(x- 8) + 3(x+ 8) = 0 now we can not factor out this equation so it is Prime ..

OpenStudy (anonymous):

that is D

OpenStudy (anonymous):

So shiraz was right?

OpenStudy (haseeb96):

yes

OpenStudy (paki):

x^2 – 5x + 24 x^2-8x+3x+24 x(x-8)+3(x+8)... .so option D...

OpenStudy (anonymous):

Ok, Thanks you guys :D

OpenStudy (anonymous):

You are welcome :)

OpenStudy (paki):

pleasure...